Paula's current salary as a systems engineer is $100,000 per year. Paula is planning to place 15% of her salary each year in a mutual fund. Paula expects a 5% salary increase each year for the next 5 years of employment. if the mutual fund will average 10% per year, what will be the amount of interest accumulated at the end of year 5?
Answer #1

Contribution to mutual fund at the end of year 1=R=100000*15%=$15000

Growth rate of contribution=growth rate of salary=g=5%

Return=i=10%

Time period of investment=n=5 years

Accumulated amount, FVg, is given by

|(1+i) - (1 +9) FVg=R* i-g

FV g = 15000 * [(1 + 0.10) – (1 + 0.05) 0.10 – 0.05

(1.610510 – 1.276281563 FV9= 15000* 0.05 = 15000*6.684569 = 100268.54

Accumulated amount=$100,268.54